find three consecutive integers whose sum is -147
Answer by mananth(16946)

let the inetegers be x, x+1,x+2
...
x+x+1+x+2=-147
3x+3=-147
-3 -3
3x=-147-3
3x=-150
/3
x= -150/3
x= -50
..
-50,-49,-48
...
m.ananth@hotmail.ca